OpenStudy (anonymous):

Please help asap will fan and medal Bo and Erica are yoga instructors. Between the two of them, they teach 47 yoga classes each week. If Erica teaches 10 fewer than twice as many as Bo, how many classes does each instructor teach per week? 16 Bo; 31 Erica 21 Bo; 26 Erica 19 Bo; 28 Erica 23 Bo; 24 Erica

OpenStudy (kamibug):

Let Bo be represented by the variable b and Erica by e. You're given that b + e = 47 and e = b - 10. This is a system of equations. You can solve it using substitution. Plug in the value of e into the first equation and solve for b. :) b + b - 10 = 47 2b - 10 = 47 2b = 57 b = 28.5 <--- Now that you know what b is, substitute it's value into either of the equations and you can solve for e. 28.5 + e = 47 e = 47 - 28.5 e = ? <---

OpenStudy (anonymous):

C.

OpenStudy (kamibug):

Right:)